OpenStudy (anonymous):

17 more than 2 times a number is the same as 5 less than 12 times the number

OpenStudy (dan815):

17+2x=12x-5

OpenStudy (dan815):

solve for x

OpenStudy (anonymous):

yes, what is the number?

OpenStudy (imstuck):

Combine the like terms to solve for x. Get all the x terms together on one side and the 17 and 5 together on the other side and solve for x.

OpenStudy (anonymous):

x=22/10?

OpenStudy (skullpatrol):

simplify please :)

OpenStudy (imstuck):

Yes!!!

OpenStudy (anonymous):

x = 2.2

OpenStudy (imstuck):

Then you could simplify by reducing.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

L=c+7d/6 , d=?